# Generalized Random Forests using Fixed-Point Trees

### v3

@@ A.2 Forest specifications
The following are specifications imposed upon the trees of a generalized random forest, and come from the forest specifications of GRF for forests with gradient-based trees (Athey et al., 2019). These specifications are used in the asymptotic analyses of GRF-FPT and describe fairly mild conditions on the tree splitting mechanism, as well as a precise sampling mechanism used to train the trees.
@@
    (Symmetric) Tree estimates are invariant to permutations of the training indices.
@@
    (Balanced/$\omega$-regular) Each split assigns at least fraction $\omega>0$ of observations from the parent to each child node.
@@
    (Randomized/random-split) The probability of splitting along any feature/dimension of the input space is at least $\pi>0$.
@@ A.3 Regularity conditions
    $\text{Var}(\rho_{i}^{*}(x)\,|\,X_{i}=x)>0,\qquad\text{for all $x\in\mathcal{X}$.}$
@@
    where $10<\pi,\omega<$ 1 are constants defined in Specifications A.2.
@@ Appendix B Derivations and Proofs
B.1 Equivalent multivariate CART criteria
Let $\rho_{i}\in\mathbb{R}^{K}$ be multivariate quantities over a parent node $P$. A standard CART split of $P$ is made by the child nodes $C_{1}$ and $C_{2}$ that minimize
@@ B.1 Equivalent multivariate CART criteria
where $\bar{\rho}_{C_{j}}\coloneqq n_{C_{j}}^{-1}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i}$ is the local prediction over child node $C_{j}$. We show that a split $\{C_{1},C_{2}\}$ minimizes the least-squares CART criterion (31) if and only if it maximizes
@@
A conventional CART least-squares criterion (31) can be written as
@@
=$=\sum_{\{i:X_{i}\in P\}}\left\lVert\rho_{i}\right\rVert^{2}-\left(n_{C_{1}}\left\lVert\bar{\rho}_{C_{1}}\right\rVert^{2}+n_{C_{2}}\left\lVert\bar{\rho}_{C_{2}}\right\rVert^{2}\right),$
where the first term does not depend on the choice of split. Therefore, a split $\{C_{1},C_{2}\}$ that minimizes (31) also maximizes (32).
B.2 Approximate tree splits via CART
The following result is a generalization to the claim that a CART split on pseudo-outcomes $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ will produce a split that maximizes the $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$-criterion, and covers the analogous result for gradient-based quantities $\rho_{i}^{\texttt{grad}}$ and $\widetilde{\Delta}^{\texttt{grad}}(C_{1},C_{2})$. Starting with some notation, let $a$ and $B$ denote vector and matrix quantities, respectively, that are independent of $\{C_{1},C_{2}\}$. Define
$\tilde{\theta}_{C_{j}}=\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}(a-B\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i}))=a+\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i},$
where $\rho_{i}$ denotes a pseudo-outcome made by a linear transformations of the parent score $\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$,
$\rho_{i}=-B\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i}).$ (33)
The following result argues that quantities $\tilde{\theta}_{C_{j}}$ made by an average of affine transformations of parent scores $\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$, lead to $\widetilde{\Delta}$-style-criteria that is maximized by a CART split on pseudo-outcomes (33).
@@ B.2 Approximate tree splits via CART
Suppose that $\{C_{1},C_{2}\}$ is a CART split on $\rho_{i}$ of the form (33) over covariates $X_{i}\in P$. Under Assumptions A.1, the child nodes $C_{1}$ and $C_{2}$ maximize
$\widetilde{\Delta}(C_{1},C_{2})=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\tilde{\theta}_{C_{1}}-\tilde{\theta}_{C_{2}}\right\rVert^{2}.$
Proof.
The scores $\psi_{\theta,\nu}(O_{i})$ satisfy subgradient conditions by assumption, and so the local solutions $(\hat{\theta}_{P},\hat{\nu}_{P})$ satisfy first-order conditions
@@
Each sum in the previous expression can be written as $\sum\rho_{i}=n_{C_{j}}(\tilde{\theta}_{C_{j}}-a)$. Hence,
@@
Therefore, letting $\bar{\rho}_{C_{j}}\coloneqq\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i}$, we have
@@
Similarly, by symmetric arguments with respect to $C_{2}$, we have
@@
By the derivation in Appendix B.1, the expression $n_{C_{1}}\left\lVert\bar{\rho}_{C_{1}}\right\rVert^{2}+n_{C_{2}}\left\lVert\bar{\rho}_{C_{2}}\right\rVert^{2}$ is maximized if $\{C_{1},C_{2}\}$ is a CART split on $\rho_{i}$ over $P$. That is, $\widetilde{\Delta}(C_{1},C_{2})$ is maximized by a split $\{C_{1},C_{2}\}$ return by CART on $\rho_{i}=-B\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$ over covariates $X_{i}\in P$, as desired. ∎
B.3 Proofs for Section 4

### v4

@@ A.2 Forest specifications
The consistency and asymptotic normality results, Theorems 4.3 and 4.4, require that the forest trained following Algorithm 2 consists of trees that satisfy a certain set of specifications. These forest specifications are precisely those imposed by Athey et al. (2019) for forests of gradient-based trees, and collectively, these specifications describe fairly mild conditions on the tree splitting mechanism, as well as specific requirements for the sampling procedure.
@@
    (Symmetric) Tree estimates are invariant to permutations of the training indices. In other words, the output of a tree does not depend on the order in which the training samples are indexed.
@@
    (Balanced/$\omega$-regular) The proportion of parent observations assigned into either child is bound below by some $\omega>0$, i.e. $n_{C_{j}}\geq\omega n_{P}$.
@@
    (Randomized/random-split) The probability of splitting along any feature/dimension of the input space is bound below by some $\pi>0$.
@@ A.3 Regularity conditions
    $\text{Var}(\rho_{i}^{*}(x)\mid X_{i}=x)>0,\qquad\text{for all $x\in\mathcal{X}$.}$
@@
    where $10<\pi,\omega<$ 1 are constants defined in forest Specifications 2 and 3.
A.4 Neyman orthogonality
To identify the underlying local parameters $(\theta^{*}(x),\nu^{*}(x))\in\mathbb{R}^{K}$ one must have a score $\psi_{\theta,\nu}(O)$ with at least $K=K_{\theta}+K_{\nu}$ components, where here we use $K_{\theta}$ and $K_{\nu}$ to denote the dimensions of the component subvectors $\theta^{*}(x)\in\mathbb{R}^{K_{\theta}}$ and $\nu^{*}(x)\in\mathbb{R}^{K_{\nu}}$. Conceptually, a score $\psi_{\theta,\nu}(O)$ can be partitioned into the components that identify the $\theta$-coordinates, denoted by $\psi_{1}$, and those that identify the $\nu$-coordinates, denoted by $\psi_{2}$, and thus the moment functions $M_{\theta,\nu}(x)$ in (36) can also be partitioned the same way:
$\psi_{\theta,\nu}(O)=\begin{bmatrix}\psi_{1}(\theta,\nu;O)\\ \psi_{2}(\theta,\nu;O)\end{bmatrix},\qquad M_{\theta,\nu}(x)=\begin{bmatrix}M_{1}(\theta,\nu;x)\\ M_{2}(\theta,\nu;x)\end{bmatrix}=\begin{bmatrix}\mathbb{E}[\psi_{1}(\theta,\nu;O)\mid X=x]\\ \mathbb{E}[\psi_{2}(\theta,\nu;O)\mid X=x]\\ \end{bmatrix}.$
The corresponding Jacobian matrix of $M_{\theta,\nu}(x)$ taken with respect to $(\theta,\nu)$ and evaluated at the truth $(\theta^{*}(x),\nu^{*}(x))$ is
$V(x)=\nabla_{(\theta,\nu)}\left.M(\theta,\nu;x)\right|_{\theta=\theta^{*}(x),\nu=\nu^{*}(x)}=\begin{bmatrix}V_{\theta\theta}(x)&V_{\theta\nu}(x)\\ V_{\nu\theta}(x)&V_{\nu\nu}(x)\end{bmatrix},$
where here the subscripts in the block expressions of $V(x)$ indicate the coordinates with which the gradient is taken, and in all cases are evaluated at the truth $(\theta^{*}(x),\nu^{*}(x))$:
$V_{\theta\theta}(x)$ $=\nabla_{\theta}\left.M_{1}(\theta,\nu;x)\right|_{\theta=\theta^{*}(x),\nu=\nu^{*}(x)},$
$V_{\theta\nu}(x)$ $=\nabla_{\nu}\left.M_{1}(\theta,\nu;x)\right|_{\theta=\theta^{*}(x),\nu=\nu^{*}(x)},$
$V_{\nu\theta}(x)$ $=\nabla_{\theta}\left.M_{2}(\theta,\nu;x)\right|_{\theta=\theta^{*}(x),\nu=\nu^{*}(x)},$
$V_{\nu\nu}(x)$ $=\nabla_{\nu}\left.M_{2}(\theta,\nu;x)\right|_{\theta=\theta^{*}(x),\nu=\nu^{*}(x)}.$
In this context, the assumption of Neyman orthogonal moment conditions is more completely labeled as Neyman orthogonality for the estimation of $\theta^{*}(x)$ with respect to the nuisance $\nu^{*}(x)$, and can be summarized as an assumption that the moment conditions for $\theta^{*}(x)$ are insensitive to first-order changes in $\nu$ around the truth $\nu^{*}(x)$ whenever $\theta=\theta^{*}(x)$. For GRF, this means that one assumes (1) satisfies $M_{1}(\theta^{*}(x),\nu^{*}(x);x)={\bf 0}$, and in other words, the partial derivatives of the moment functions for $\theta^{*}(x)$ with respect to $\nu$ are zero at $(\theta^{*}(x),\nu^{*}(x))$:
$V_{\theta\nu}(x)={\bf 0}.$
A.5 Example: Neyman orthogonality for VCM and HTE
Consider the VCM/HTE model with data $(Y_{i},W_{i},X_{i})$ related according to
$\mathbb{E}[Y_{i}\mid X_{i}=x]=\nu^{*}(x)+W_{i}^{\top}\theta^{*}(x),$
such that, as discussed in Section 3.2, the score function $\psi_{\theta,\nu}$ that identifies the underlying $(\theta^{*}(x),\nu^{*}(x))$ is
$\psi_{\theta,\nu}(Y_{i},W_{i})\coloneqq\begin{bmatrix}(Y_{i}-W_{i}^{\top}\theta-\nu)W_{i}\\ Y_{i}-W_{i}^{\top}\theta-\nu\end{bmatrix},$
and the corresponding local Jacobian $V(x)$ has block form
$V(x)=-\mathbb{E}\left[\begin{bmatrix}W_{i}W_{i}^{\top}&W_{i}^{\top}\\ W_{i}&1\end{bmatrix}\;\middle|\;X_{i}=x\right]=-\begin{bmatrix}\mathbb{E}[W_{i}W_{i}^{\top}\mid X_{i}=x]&\mathbb{E}[W_{i}^{\top}\mid X_{i}=x]\\ \mathbb{E}[W_{i}\mid X_{i}=x]&1\end{bmatrix}.$
Therefore, for Neyman orthogonality to hold one requires that $\mathbb{E}[W_{i}\mid X_{i}=x]={\bf 0}$.
@@ Appendix B Derivations and Proofs
B.1 Proofs for Section 3.4
B.1.1 Multivariate CART criteria
Let $\rho_{i}\in\mathbb{R}^{K}$ be vector-valued responses associated with covariates $X_{i}\in P$. A standard CART split $(C_{1},C_{2})$ of $P$ minimizes the conventional least-squares criterion:
@@ B.1.1 Multivariate CART criteria
where $\bar{\rho}_{C_{j}}\coloneqq n_{C_{j}}^{-1}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i}$ is the local prediction over child node $C_{j}$. We verify that a split $(C_{1},C_{2})$ minimizes (38) if and only if it maximizes
@@
Therefore, the least-squares criterion CART (38) is equivalently written as
@@
=$=\sum_{\{i:X_{i}\in P\}}\left\lVert\rho_{i}\right\rVert^{2}-\left(n_{C_{1}}\left\lVert\bar{\rho}_{C_{1}}\right\rVert^{2}+n_{C_{2}}\left\lVert\bar{\rho}_{C_{2}}\right\rVert^{2}\right).$
The first term does not depend on the choice of split, and therefore the split that minimizes (38) is equivalent to the split that maximizes (39). ∎
B.1.2 Splits via CART on pseudo-outcomes
The following result is a generalization to the claim made in Section 3.4 that a CART split on pseudo-outcomes $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ will produce a split that maximizes the $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$-criterion, and is sufficiently general to cover gradient-based pseudo-outcomes $\rho_{i}^{\texttt{grad}}$ and the corresponding $\widetilde{\Delta}^{\texttt{grad}}$-criterion.
@@ B.1.2 Splits via CART on pseudo-outcomes
Suppose we can write
$\tilde{\theta}_{C_{j}}=a+\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i},\qquad\rho_{i}=-B\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i}),$ (40)
where $a$ and $B$ denote appropriately sized vectors and matrices whose values do not depend on the candidate child node $C_{j}$. Under Assumptions A.1, the split $(C_{1},C_{2})$ that maximizes
$\widetilde{\Delta}(C_{1},C_{2})=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\tilde{\theta}_{C_{1}}-\tilde{\theta}_{C_{2}}\right\rVert^{2},$
is exactly the split chosen by CART for vector-valued responses $\rho_{i}$ fit over covariates $X_{i}\in P$.
Proof of Lemma B.1.
The scores $\psi_{\theta,\nu}(O_{i})$ satisfy subgradient conditions by Assumption 6, and therefore the parent solutions $(\hat{\theta}_{P},\hat{\nu}_{P})$ satisfy the first-order conditions
@@
Each sum in the previous expression is equivalently written as $\sum\rho_{i}=n_{C_{j}}(\tilde{\theta}_{C_{j}}-a)$. Hence,
@@
Writing $\bar{\rho}_{C_{j}}\coloneqq\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i}$, one has:
@@
Applying analogous arguments with respect to $C_{2}$, one has the symmetric result:
@@
Based on the arguments in Appendix B.1.1, a split $(C_{1},C_{2})$ maximizes $n_{C_{1}}\left\lVert\bar{\rho}_{C_{1}}\right\rVert^{2}+n_{C_{2}}\left\lVert\bar{\rho}_{C_{2}}\right\rVert^{2}$ if and only if it is a CART split performed on the $\rho_{i}$ over $P$. That is, $\widetilde{\Delta}(C_{1},C_{2})$ is precisely maximized by a single CART split on $\rho_{i}=-B\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i})$ fit over covariates $X_{i}\in P$, as desired. ∎
B.1.3 Scale invariance of CART splits
Lemma B.2 (Argmax equivalence of FPT criteria).
The optimal split identified by CART on pseudo-outcomes $\rho_{i}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ of the form (23) does not depend on the scale factor $\eta$, for any $\eta\neq 0$.
Proof of Lemma B.2.
Denote by $\rho_{i}^{(\eta)}$ FPT pseudo-outcomes based on an arbitrary scale factor $\eta\neq 0$ of the form (23):
$\rho_{i}^{(\eta)}\coloneqq-\eta\xi^{\top}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i}),$ (41)
and let $\overline{\psi}_{C_{j}}$ denote the child-leaf average score evaluated at the parent solution $(\hat{\theta}_{P},\hat{\nu}_{P})$:
$\overline{\psi}_{C_{j}}\coloneqq\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\psi_{\hat{\theta}_{P},\hat{\nu}_{P}}(O_{i}),$
such that the corresponding child-leaf pseudo-outcome averages $\bar{\rho}_{C_{j}}^{~{}(\eta)}\coloneqq\frac{1}{n_{C_{j}}}\sum_{\{i:X_{i}\in C_{j}\}}\rho_{i}^{(\eta)}$ are equivalently written as
$\bar{\rho}_{C_{j}}^{~{}(\eta)}=-\eta\xi^{\top}~{}\overline{\psi}_{C_{j}}$
Let $\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ denote the FPT criterion of the form (25) based on pseudo-outcomes (41):
$\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\bar{\rho}_{C_{1}}^{~{}(\eta)}-\bar{\rho}_{C_{j}}^{~{}(\eta)}\right\rVert^{2}=\frac{n_{C_{1}}n_{C_{2}}}{n_{P}^{2}}\left\lVert\eta\xi^{\top}(\overline{\psi}_{C_{1}}-\overline{\psi}_{C_{2}})\right\rVert^{2}.$
One has:
$\left\lVert\eta\xi^{\top}(\overline{\psi}_{C_{1}}-\overline{\psi}_{C_{2}})\right\rVert^{2}=\eta^{2}\left\lVert\xi^{\top}(\overline{\psi}_{C_{1}}-\overline{\psi}_{C_{2}})\right\rVert^{2},$
and hence the $\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$-criteria obey the scaling relation:
$\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})=\eta^{2}\cdot\widetilde{\Delta}_{1}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2}),$ (42)
where $\widetilde{\Delta}_{1}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$ denotes the FPT criterion induced by pseudo-outcomes $\rho_{i}^{(1)}$ based on unit scale factor $\eta=1$. The relation (42) implies that any nonzero split-independent rescaling $\rho_{i}^{(\eta)}=\eta\rho_{i}^{(1)}$ will induce a splitting criterion $\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$ with the same maximizer as $\widetilde{\Delta}_{1}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})$:
$\operatorname*{arg\,max}_{(C_{1},C_{2})}~{}\left\{\widetilde{\Delta}_{\eta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})\right\}=\operatorname*{arg\,max}_{(C_{1},C_{2})}~{}\left\{\eta^{2}\cdot\widetilde{\Delta}_{1}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})\right\}=\operatorname*{arg\,max}_{(C_{1},C_{2})}~{}\left\{\widetilde{\Delta}_{1}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}(C_{1},C_{2})\right\}.$
Intuitively, a CART split is chosen by ranking the criterion values among the candidate splits and selecting the maximizing split $(C_{1},C_{2})$. Therefore, the FPT splitting mechanism is unaffected by the scale factor $\eta$ used to specify fixed-point pseudo-outcomes (23). The absolute scale of the $\widetilde{\Delta}^{\texttt{\color[rgb]{0.5,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0,0}\pgfsys@color@cmyk@stroke{0}{0.85}{0.85}{0.50}\pgfsys@color@cmyk@fill{0}{0.85}{0.85}{0.50}FPT}}$-criterion does not matter when searching for the optimal split, and only the criterion rankings across the candidate splits determine the final partition. ∎
B.2 Proofs for Section 4